# Closed geodesics in dilation surfaces

### v2

@@ 3.6. Delaunay limits
As mentioned above, we will argue by induction on the complexity of the surface $\ \Sigma$ (see Definition 2.4). It is easy to deal with the case of smallest complexity, the case of flat tori formed by a pair of triangles. The induction assumption is that any dilation surface of complexity lower than $k$ carries cylinders in a dense set of directions. We want to show that it is still the case for surfaces of complexity $k+1$.
It follows from Lemma 2.8 that for a dilation surface $\Sigma$ without boundary, an open set of $\mathbb{RP}^{1}$ that does not contain any direction of saddle connection contains the direction of a closed geodesic. Therefore, it remains to prove that any direction $\theta\in\mathbb{RP}^{1}$ that is approached by directions of saddle connections is also approached by directions of closed geodesics. In other words, any open subset of $\mathbb{RP}^{1}$ containing the direction of a saddle connection should also contain the direction of a closed geodesic. Up to the action of an element of $\mathrm{SL}_{2}({\mathbb{R}})$, we can assume that $\Sigma$ contains a vertical saddle connection $\gamma$ and $U$ is an open subset of $\mathbb{RP}^{1}$ contains the vertical direction.
To rely on the induction assumption, we will use the Teichmüller flow $g_{t}$ that contracts the vertical direction, and therefore the saddle connection $\gamma$, and that expands the horizontal direction. We have seen in Section 3.3 that one can extract a sequence of times $t_{n}\to+\infty$ such that the sequence $(\Sigma_{n})_{n\in\mathbb{N}}:=(g_{t_{n}}\Sigma)_{n\in\mathbb{N}}$ is Delaunay-convergent. We first rule out the case in which $(\Sigma_{n})_{n\in\mathbb{N}}$ Delaunay-converges toward a surface $\Sigma_{\infty}$ of same complexity (see Section 3.6 for a precise definition of Delaunay limits).
@@
Before entering the proof of the above proposition, let us mention that there is a class of dilation surfaces, called quasi-Hopf surfaces (see [7] for details), having a vertical saddle connection and whose Teichmüller orbit is periodic. These surfaces decompose into disjoint dilation cylinders whose boundary saddle connections are either horizontal or vertical.
@@
The limit surface $\Sigma_{\infty}$ carries at least one vertical saddle connection.
@@
On a dilation surface, vertical saddle connections have a top and a bottom endpoint. For any vertical saddle connection $\gamma$, we denote by $\mathcal{R}(\gamma)$ the vertical leaf satisfying the following properties:
@@
  • at $M$, $\mathcal{R}(\gamma)$ and $\gamma$ form an angular sector of amplitude $\pi$ contained in the right half-plane ($\mathcal{R}(\gamma)$ is obtained from $\gamma$ by turning counter-clockwise around $M$ by an angle of $\pi$).
@@
We argue by contradiction, assuming that $\Sigma_{\infty}$ contains a vertical saddle connection $\gamma$ such that $\mathcal{R}(\gamma)$ is not a saddle connection. We denote by $s_{\mathrm{top}}$ the top singularity of $\gamma$ and choose a continuous parametrization $\Gamma(t)$ of the leaf $\mathcal{R}(\gamma)$ with $\Gamma(0)=s_{\mathrm{top}}$, see Figure 12. For $t$ small enough we define $\mathcal{D}_{t}$ the unique immersed disk whose center is $\Gamma(t)$ and whose radius is given by the segment $[s_{\mathrm{top}},\Gamma(t)]$. Note that for any $t>t^{\prime}$ we have $\mathcal{D}_{t^{\prime}}\subset\mathcal{D}_{t}$. We then define the "closest" singularity $s_{\mathrm{top}}^{\min}$ to $\gamma_{\infty}$ (on the right side of $\gamma_{\infty}$) as the first singularity encountered when considering the increasing sequence of disks $(\mathcal{D}_{t})_{t\geq 0}$. Note that such a sequence must actually encounter a singularity because of Theorem 2.11.
@@ 6.2. Cyclic maximal domains of type 2
The proof is based on Proposition 2.9. This proposition asserts that, in the case of a dilation surface with boundary, either a given open set of directions contains a cylinder or a set of trajectories having these directions, a pencil to be precise, must leave across a boundary component of the dilation surface. This proposition then shows that we can concentrate on the case where each trajectory of a Delaunay piece (see Definition 3.12) leaves it by hitting the boundary. It can do it by crossing either a long edge or a short one. We will actually rule out the short edge case in Sections 7.1 and 7.2. Indeed, these boundaries are by definition very small compared to the long edges and it will be unlucky to leave the piece through such a short side. The trajectories of the pencil will then have to leave the Delaunay piece through a long side and then enter a new Delaunay piece through a short side (as by construction Delaunay pieces are glued to one another short side to long side). If the pencil does not enter in a cylinder one can repeat the argument to get a sequence of Delaunay pieces such that the pencil enters them by short side and leave them by long sides. As there are only finitely many boundary components, such a pencil will cross twice a given edge. The first return map on such an edge is a very dilating mapping as going from short sides to long sides induces a huge contraction. This concludes as contracting mappings have periodic orbits.

### v3

@@ 3.6. Delaunay limits
As mentioned above, we will argue by induction on the complexity of the closed surface $\Sigma$ (see Definition 2.4). It is easy to deal with the case of smallest complexity, the case of flat tori formed by a pair of triangles. The induction assumption is that any closed dilation surface of complexity lower than $k$ carries cylinders in a dense set of directions. We want to show that it is still the case for surfaces of complexity $k+1$.
It follows from Lemma 2.8 that for a dilation surface $\Sigma$ without boundary, an open set of $\mathbb{RP}^{1}$ that does not contain any direction of saddle connection contains the direction of a closed geodesic. Therefore, it remains to prove that any direction $\theta\in\mathbb{RP}^{1}$ that is approached by directions of saddle connections is also approached by directions of closed geodesics. In other words, any open subset of $\mathbb{RP}^{1}$ containing the direction of a saddle connection should also contain the direction of a closed geodesic. Up to the action of an element of $\mathrm{SL}_{2}({\mathbb{R}})$, we can assume that $\Sigma$ contains a vertical saddle connection $\gamma$ and $U$ is an open subset of $\mathbb{RP}^{1}$ containing the vertical direction.
To rely on the induction assumption, we will use the Teichmüller flow $g_{t}$ that contracts the vertical direction, and therefore the saddle connection $\gamma$, and that expands the horizontal direction. We have seen in Section 3.3 that one can extract a sequence of times $t_{n}\to+\infty$ such that the sequence $(\Sigma_{n})_{n\in\mathbb{N}}:=(g_{t_{n}}\Sigma)_{n\in\mathbb{N}}$ is Delaunay-convergent. We are doing induction on closed surfaces and this is why the case where $\Sigma_{\infty}$ has a boundary will need to be handled separately. We first rule out the case in which $(\Sigma_{n})_{n\in\mathbb{N}}$ Delaunay-converges toward a closed surface $\Sigma_{\infty}$ of same complexity (see Section 3.6 for a precise definition of Delaunay limits).
@@
Before entering the proof of the above proposition, let us mention that there is a class of dilation surfaces, called quasi-Hopf surfaces (see [7] for details), having a vertical saddle connection and whose Teichmüller orbit is periodic. These surfaces decompose into disjoint dilation cylinders whose boundary saddle connections are either horizontal or vertical. We cannot extract from the Teichmüller orbit of these surfaces a sequence that Delaunay-converges to a surface of smaller complexity.
@@
Let $\Sigma$ be a closed dilation surface that carries a vertical saddle connection and a sequence of times $t_{n}\to+\infty$ such that $(g_{t_{n}}\Sigma)_{n\in\mathbb{N}}$ Delaunay-converges toward a surface $\Sigma_{\infty}$ having the same complexity as $\Sigma$. Then the limit surface $\Sigma_{\infty}$ carries at least one vertical saddle connection.
@@
On a dilation surface, vertical saddle connections have a top and a bottom endpoint. For any vertical saddle connection $\gamma$, we denote by $\mathcal{R}(\gamma)$ the vertical ray satisfying the following properties:
@@
  • at $M$, $\mathcal{R}(\gamma)$ and $\gamma$ form an angular sector of amplitude $\pi$ contained in the right half-plane (by convention). In other words, $\mathcal{R}(\gamma)$ is obtained from $\gamma$ by turning counter-clockwise around $M$ by an angle of $\pi$.
@@
We argue by contradiction, assuming that $\Sigma_{\infty}$ contains a vertical saddle connection $\gamma$ such that $\mathcal{R}(\gamma)$ is not a saddle connection. We denote by $s_{\mathrm{top}}$ the top singularity of $\gamma$ and choose a continuous parametrization $\Gamma(t)$ of the ray $\mathcal{R}(\gamma)$ with $\Gamma(0)=s_{\mathrm{top}}$, see Figure 12. For $t$ small enough we define $\mathcal{D}_{t}$ the unique immersed disk whose center is $\Gamma(t)$ and whose radius is given by the segment $[s_{\mathrm{top}},\Gamma(t)]$. Note that for any $t>t^{\prime}$ we have $\mathcal{D}_{t^{\prime}}\subset\mathcal{D}_{t}$. We then define the "closest" singularity $s_{\mathrm{top}}^{\min}$ to $\gamma_{\infty}$ as the first singularity encountered when considering the increasing sequence of disks $(\mathcal{D}_{t})_{t\geq 0}$. Note that such a sequence must actually encounter a singularity because of Theorem 2.11.
@@ 6.2. Cyclic maximal domains of type 2
The proof is based on Proposition 2.9. This proposition asserts that, in the case of a dilation surface with boundary, either a given open set of directions contains a cylinder or a set of trajectories having these directions, a pencil to be precise, must leave across a boundary component of the dilation surface. This proposition then shows that we can concentrate on the case where each trajectory of a Delaunay piece (see Definition 3.12) leaves it by hitting the boundary. It can do it by crossing either a long edge or a short one. We will actually rule out the short edge case in Sections 7.1 and 7.2. Indeed, these boundaries are by definition very small compared to the long edges and it will be unlucky to leave the piece through such a short side. The trajectories of the pencil will then have to leave the Delaunay piece through a long side and then enter a new Delaunay piece through a short side (as by construction Delaunay pieces are glued to one another short side to long side, see Lemma 3.13). If the pencil does not enter in a cylinder one can repeat the argument to get a sequence of Delaunay pieces such that the pencil enters them by short side and leave them by long sides. As there are only finitely many boundary components, such a pencil will cross twice a given edge. The first return map on such an edge is a very dilating mapping as going from short sides to long sides induces a huge contraction. This concludes as contracting mappings have periodic orbits.